OpenStudy (anonymous):

Help Again I been having a hard time trying to solve for this! find the area under one arch of the trochoid: x=6(theta)−4 sin(theta) y=6−4 cos(theta)

OpenStudy (anonymous):

@mathmath333

OpenStudy (dan815):

ah okay!

OpenStudy (dan815):

what have you tried so far

OpenStudy (anonymous):

ok so I tried: Because I know A= \[\int\limits_{\alpha}^{\beta}(g(x))(f'(x)) so then: ok so I tried: Because I know A= \[\int\limits_{\pi}^{\2pi}(6(theta)-4sin(theta)(-4sin(theta))

OpenStudy (anonymous):

yeah I am not sure there

OpenStudy (dan815):